# SS4Rec: Continuous-Time Sequential Recommendation with State Space Models

### v2

@@ Abstract
Index Terms:
State Space Model, Continuous-Time Modeling, Sequential Recommendation.
@@ I Introduction
Recommender systems are widely employed in social media and e-commerce to alleviate information overload [1]. The core of recommender systems is accurately predicting user interests in order to suggest appropriate items for users from a large candidate pool [2]. Notably, user interests are always changing continuously over time, and this evolution is reflected in their historical interaction sequence. As a result, sequential recommender systems have been developed to offer personalized online services by capturing the users’ evolving interests through their interaction sequences.
@@
Earlier sequential recommendation approaches utilized Markov chains to capture sequential patterns by modeling the transition probabilities between items [3, 4]. With the advent of deep neural networks, a series of related methods emerged, including those based on recurrent neural networks (RNNs) [5, 6, 7], convolutional neural networks (CNNs) [8], and Transformer models [9, 10, 11]. Although these approaches achieve promising performance, they are unable to effectively capture long-range dependencies and suffer from the inference inefficiency problem. Additionally, these models are inherently discrete and designed for dealing with discrete systems, however, user interests evolve continuously and vary over time, which requires recommender systems to adapt to the user’s current state as well as the dynamic interest evolution.
@@
Recently, state space models (SSMs) have become an alternative to neural network-based models due to their effectiveness in modeling continuous dynamic systems. S4 [12] and S5 [13] utilize HiPPO matrix initialization and parallel scanning for long-range reasoning and fast inferring, respectively. The latest SSM, Mamba (i.e., S6) [14], employs an input-dependent selective state space model to retain input-relevant information while filtering out irrelevant information. This improves its sequential modeling capability and inspires researchers to explore the potential of SSMs for sequential recommendation. Mamba4Rec [15] and RecMamba [16] leverages Mamba to capture sequential dependencies from historical interactions, modeling complex relationships between users and items while partially addressing the trade-off between recommendation performance and inference efficiency.
@@
Despite the success of SSMs in efficient sequential recommendations for long sequences and large-scale datasets, these methods are still under the assumption of uniform time intervals, as Mamba was originally designed for NLP tasks. Therefore, existing SSM-based approaches focus on capturing relation-aware information but overlook the impact of varying temporal dynamics in user interest evolution. In reality, the observed interactions are often recorded as discrete time series data with irregular time intervals [17], which play a crucial role in maintaining the underlying continuity. Although some methods [18, 19, 20, 21] introduce time-dependent embeddings or kernel functions to model temporal information, they still struggle to overcome the assumption of uniform time intervals. This limitation disrupts the underlying continuity of the recommender system, impairing its ability to capture users’ evolving interests and time-specific preferences (e.g., morning vs. evening).
@@ II-A Sequential Recommendation
Sequential recommendation has been extensively studied, with early approaches primarily relying on Markov chain. For example, FPMC (Factorizing Personalized Markov Chains) [3] combines the factorization machine and Markov chain techniques to construct a personalized item-item relationship transition matrix for each user, enabling the next-item prediction based on users’ most recent interactions.
In recent years, deep neural networks (DNNs) have gained significant attention due to their strong expressive capabilities, leading to the emergence of DNN-based models in the field of sequential recommendation. GRU4Rec [6] utilizes Gate Recurrent Unit (GRU) [22] network with ranking loss to capture the user interests within the sequence. NARM [23] applies GRU to extract the users’ general interests and intentions through a global and local encoder, respectively. SASRec [10] employs the Transformer-based architecture to model user interaction sequences, capturing long-term dependencies while adaptively adjusting attention weights of items in the sequence, thereby focusing on important information. Bert4Rec [11] introduces a Cloze task and utilizes a bidirectional attention mechanism to learn contextual information from the interaction sequences. DCRec [24] and GCL4SR [25] leverage self-supervised learning techniques to capture complex relationships within sequential data and enhance recommendation performance. All the aforementioned approaches are devoted to modeling sequential dependencies in historical user behaviors but ignore the interaction timestamps, limiting their ability to capture the dynamic evolution of user interests.
@@ II-B Time-Aware Recommendation
Several time-aware sequential recommendation approaches have been proposed to incorporate temporal information in modeling user preferences. TiSASRec [18] adopts a time-aware self-attentive model to learn the weights of different items in a sequence, integrating both absolute positions and relative time intervals between items to predict future items. TGSRec [26] combines sequential patterns and temporal collaborative signals by introducing a Temporal Collaborative Transformer (TCT) layer, capturing both user-item interactions and temporal dynamics for continuous-time recommendation. TAT4SRec [20] processes item sequences and timestamp sequences separately, using a window function to maintain continuity in timestamps. TASER [19] introduces a relative temporal attention network to capture the pairwise dependencies using absolute temporal vectors. The work of TiCoSeRec [27] demonstrates that uniformly distributed time intervals can improve performance. Different from these time-discretized approaches, we explore the continuous-time recommender systems, aiming to extract dynamic evolving information underlying continuous systems.
@@ II-C State Space Models
State space models have emerged as a promising class of architectures, demonstrating exceptional performance in sequence modeling tasks [28]. The first structured state space sequence model, S4 [12], addresses the challenge of capturing long-range dependencies through the state matrix initialization based on HiPPO framework [29]. Different from the single-input, single-output scheme of S4, S5 [13] adopts multi-input, multi-output (MIMO) SSM. By utilizing a diagonal state matrix, S5 achieves linear computational complexity and time-varying parameterization via parallel scans. The latest work of SSM, Mamba [14], incorporates a selection mechanism to filter irrelevant information while retaining relevant information, and designs a hardware-aware algorithm for both efficient training and inference. Mamba2 [30] connects attention variants with selective SSM directly for large-scale language modeling. Jamba [31] is a hybrid architecture for large-scale modeling that combines Mamba with Transformer and mixture-of-experts module to take full advantage of each architecture.
Recent works have introduced Mamba into recommender systems due to its impressive performance in sequential tasks. Mamba4Rec [15] is the first to apply the Mamba block for sequential recommendation with improving performance and fast inference. RecMamba [16] replaces the Transformer block in SASRec with Mamba for lifelong sequences. MaTrRec [32] combines Mamba with Transformer, leveraging their strengths of capturing both long-range and short-range dependencies. SSD4Rec [33] introduces a bi-directional structured state-space duality block to capture contextual information, similar to Bert4Rec. Despite these advancements, the challenge of time-specific personalized recommendations with irregular intervals remains largely unexplored. By formulating this problem as a continuous time-varying system and adopting hybrid state space models to capture both temporal and sequential information of user interactions, our work sheds light on the continuous-time sequential recommendation.
@@ III-C State Space Model
where $\mathbf{A}\in\mathbb{R}^{P\times P}$, $\mathbf{B}\in\mathbb{R}^{P\times D}$, $\mathbf{C}\in\mathbb{R}^{D\times P}$ are learnable parameter matrices. To employ it on a discrete input interaction sequence $[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}]$ instead of the continuous function $x(t)$, continuous-time SSM in (1) and (2) can be discretized by a constant step size $\mathbf{\Delta}$ through the zero-order hold rule:
@@ III-F Interpretation of $\mathbf{\Delta}$
$&\text{Time-Aware SSM: }&\mathbf{\Delta}\leftarrow f_{\theta}(\mathbf{\Delta T})\\ &\text{Relation-Aware SSM: }&\mathbf{\Delta}\leftarrow f_{\theta}(\mathbf{X})\\ \end{aligned}\right.$ (10)
@@ III-H Continuous-Time Prediction
where $v_{next\_time}$ is the interaction item at time $next\_time$, $\mathbf{o}\in\mathbb{R}^{D}$ denotes the last item embedding in the output embedding of the last SS4Rec block, and $\mathbf{x_{i}}\in\mathbb{R}^{D}$ denotes the candidate item embedding. In our model, the learned representation contains the information of the next interaction time, i.e., the time when an item needs to be recommended for the user. Thereby, SS4Rec is capable of selecting the most appropriate items at the optimal time for personalized recommendations.
@@ IV-A1 Baselines
  • GRU4Rec [6]: it employs Gated Recurrent Unit (GRU) to model user sequences and encode user interests into a final hidden state for session-based recommendation, where each user’s sequence of clicks is treated as a session.
  • NARM [23]: a GRU-based method adopts an encoder-decoder architecture to capture both user’s sequential actions and primary intent within a session.
@@
  • SASRec [10]: this is a classical sequential recommendation method that leverages the Transformer’s self-attention mechanism to process the input sequence.
  • BERT4Rec [11]: it utilizes bidirectional self-attention to model user behavior sequences, allowing each item to incorporated contextual information from both preceding and subsequent interactions in the user’s history.
  • TiSASRec [18]: this approach introduces discrete time interval embedding in SASRec, enabling it to effectively capture temporal information.
@@
  • Mamba4Rec [15]: the first Mamba-based sequential recommender, which mainly replaces the self-attention mechanism with the Mamba layer, in contrast to SASRec.
  • RecMamba [16]: this is a simplified model that uses only the mamba layer, which can be referred to as S6-only model in the Ablation Study.
@@ IV-A2 Datasets
  • Amazon [34]: a series of datasets are gathered from the Amazon e-commerce platform, comprising a substantial volume of product reviews extracted from the Amazon website. We extract the categories ”video games” and ”sports” for subsequent experiments.
  • Movielens [35]: a widely used dataset collected from the Movielens website records user ratings of movies seen. The version (MovieLens-1M) used in our experiments contains 1 million user ratings.
  • KuaiRec [36]: a real-world dataset is collected from the recommendation logs of the video-sharing mobile app Kuaishou. It contains two interaction matrices, a large matrix and a small matrix, where the small matrix, referred to as the first dataset, contains a fully observed user-item interaction matrix. In this paper, KuaiRec(Large) refers to the large interaction matrix, while KuaiRec(Small) denotes the small one.
@@ IV-A3 Implementation Details
To assess the efficacy of our sequential recommendation model, we follow the standard evaluation protocol used in this field. Specifically, we apply the leave-one-out strategy for testing within the recbole1.0 code framework [37]. Hit Ratio@K(HR@K), Normalized Discounted Cumulative Gain@K(NDCG@K), and Mean Reciprocal Rank@K(MRR@K) are utilized to evaluate the recommendation performance, with $K=10$ for all metrics. Higher values for these metrics signify better ranking accuracy, which indicates a more effective recommendation model. The specific hyperparameter settings are depicted in Table II, and all experiments are conducted on a single A800 80GB GPU.
@@ IV-C Time-Dependent Prediction
1:$itemSet\leftarrow[0,1,\ldots,99]$
2:$userSet\leftarrow[0,1,\ldots,99]$
3:$SeqLen\leftarrow 100$
4:$data\leftarrow[]$

### v3

@@ I Introduction
Recommender systems are widely employed in social media and e-commerce to alleviate information overload [wu2022survey, xiao2025pt4rec]. The core of recommender systems is accurately predicting user interests in order to suggest appropriate items for users from a large candidate pool [wang2019sequential]. Notably, user interests are always changing continuously over time, and this evolution is reflected in their historical interaction sequence. As a result, sequential recommender systems have been developed to offer personalized online services by capturing the users’ evolving interests through their interaction sequences.
@@
Earlier sequential recommendation approaches utilized Markov chains to capture sequential patterns by modeling the transition probabilities between items [rendle2010factorizing, he2016fusing]. With the advent of deep neural networks, a series of related methods emerged, including those based on recurrent neural networks (RNNs) [Medsker, hidasi2015session, tan2016improved], convolutional neural networks (CNNs) [tang2018personalized], and Transformer models [vaswani2017attention, kang2018self, sun2019bert4rec]. Although these approaches achieve promising performance, they are unable to effectively capture long-range dependencies and suffer from the inference inefficiency problem. Additionally, these models are inherently discrete and designed for dealing with discrete systems, however, user interests evolve continuously and vary over time, which requires recommender systems to adapt to the user’s current state as well as the dynamic interest evolution.
@@
Recently, state space models (SSMs) have become an alternative to neural network-based models due to their effectiveness in modeling continuous dynamic systems. S4 [gu2021efficiently] and S5 [smith2022simplified] utilize HiPPO matrix initialization and parallel scanning for long-range reasoning and fast inferring, respectively. The latest SSM, Mamba (i.e., S6) [gu2023mamba], employs an input-dependent selective state space model to retain input-relevant information while filtering out irrelevant information. This improves its sequential modeling capability and inspires researchers to explore the potential of SSMs for sequential recommendation. Mamba4Rec [liu2024mamba4rec] and RecMamba [yang2024uncovering] leverages Mamba to capture sequential dependencies from historical interactions, modeling complex relationships between users and items while partially addressing the trade-off between recommendation performance and inference efficiency.
@@
Despite the success of SSMs in efficient sequential recommendations for long sequences and large-scale datasets, these methods are still under the assumption of uniform time intervals, as Mamba was originally designed for NLP tasks. Therefore, existing SSM-based approaches focus on capturing relation-aware information but overlook the impact of varying temporal dynamics in user interest evolution. In reality, the observed interactions are often recorded as discrete time series data with irregular time intervals [chen2024contiformer], which play a crucial role in maintaining the underlying continuity. Although some methods [li2020time, ye2020time, zhang2023time, wu2020deja] introduce time-dependent embeddings or kernel functions to model temporal information, they still struggle to overcome the assumption of uniform time intervals. This limitation disrupts the underlying continuity of the recommender system, impairing its ability to capture users’ evolving interests and time-specific preferences (e.g., morning vs. evening).
@@ II-A Sequential Recommendation
Sequential recommendation has been extensively studied, with early approaches primarily relying on Markov chain. For example, FPMC (Factorizing Personalized Markov Chains) [rendle2010factorizing] combines the factorization machine and Markov chain techniques to construct a personalized item-item relationship transition matrix for each user, enabling the next-item prediction based on users’ most recent interactions.
In recent years, deep neural networks (DNNs) have gained significant attention due to their strong expressive capabilities, leading to the emergence of DNN-based models in the field of sequential recommendation. GRU4Rec [hidasi2015session] utilizes Gate Recurrent Unit (GRU) [Cho] network with ranking loss to capture the user interests within the sequence. NARM [li2017neural] applies GRU to extract the users’ general interests and intentions through a global and local encoder, respectively. SASRec [kang2018self] employs the Transformer-based architecture to model user interaction sequences, capturing long-term dependencies while adaptively adjusting attention weights of items in the sequence, thereby focusing on important information. Bert4Rec [sun2019bert4rec] introduces a Cloze task and utilizes a bidirectional attention mechanism to learn contextual information from the interaction sequences. DCRec [Yang2023DCRec] and GCL4SR [zhang2022GCL4SR] leverage self-supervised learning techniques to capture complex relationships within sequential data and enhance recommendation performance. All the aforementioned approaches are devoted to modeling sequential dependencies in historical user behaviors but ignore the interaction timestamps, limiting their ability to capture the dynamic evolution of user interests.
@@ II-B Time-Aware Recommendation
Several time-aware sequential recommendation approaches have been proposed to incorporate temporal information in modeling user preferences. TiSASRec [li2020time] adopts a time-aware self-attentive model to learn the weights of different items in a sequence, integrating both absolute positions and relative time intervals between items to predict future items. TGSRec [TGSRec] combines sequential patterns and temporal collaborative signals by introducing a Temporal Collaborative Transformer (TCT) layer, capturing both user-item interactions and temporal dynamics for continuous-time recommendation. TAT4SRec [zhang2023time] processes item sequences and timestamp sequences separately, using a window function to maintain continuity in timestamps. TASER [ye2020time] introduces a relative temporal attention network to capture the pairwise dependencies using absolute temporal vectors. The work of TiCoSeRec [dang2023uniform] demonstrates that uniformly distributed time intervals can improve performance. Different from these time-discretized approaches, we explore the continuous-time recommender systems, aiming to extract dynamic evolving information underlying continuous systems.
@@ II-C State Space Models
State space models have emerged as a promising class of architectures, demonstrating exceptional performance in sequence modeling tasks [xu2024integrating]. The first structured state space sequence model, S4 [gu2021efficiently], addresses the challenge of capturing long-range dependencies through the state matrix initialization based on HiPPO framework [gu2020hippo]. Different from the single-input, single-output scheme of S4, S5 [smith2022simplified] adopts multi-input, multi-output (MIMO) SSM. By utilizing a diagonal state matrix, S5 achieves linear computational complexity and time-varying parameterization via parallel scans. The latest work of SSM, Mamba [gu2023mamba], incorporates a selection mechanism to filter irrelevant information while retaining relevant information, and designs a hardware-aware algorithm for both efficient training and inference. Mamba2 [dao2024transformers] connects attention variants with selective SSM directly for large-scale language modeling. Jamba [lieber2024jamba] is a hybrid architecture for large-scale modeling that combines Mamba with Transformer and mixture-of-experts module to take full advantage of each architecture.
Recent works have introduced Mamba into recommender systems due to its impressive performance in sequential tasks. Mamba4Rec [liu2024mamba4rec] is the first to apply the Mamba block for sequential recommendation with improving performance and fast inference. RecMamba [yang2024uncovering] replaces the Transformer block in SASRec with Mamba for lifelong sequences. MaTrRec [zhang2024matrrec] combines Mamba with Transformer, leveraging their strengths of capturing both long-range and short-range dependencies. SSD4Rec [qu2025ssd4rec] introduces a bi-directional structured state-space duality block to capture contextual information, similar to Bert4Rec. Despite these advancements, the challenge of time-specific personalized recommendations with irregular intervals remains largely unexplored. By formulating this problem as a continuous time-varying system and adopting hybrid state space models to capture both temporal and sequential information of user interactions, our work sheds light on the continuous-time sequential recommendation.
@@ III-C State Space Model
where $\mathbf{A}\in\mathbb{R}^{P\times P}$, $\mathbf{B}\in\mathbb{R}^{P\times D}$, $\mathbf{C}\in\mathbb{R}^{D\times P}$ are learnable parameter matrices. To employ it on a discrete input interaction sequence $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}]$ instead of the continuous function $x(t)$, continuous-time SSM in (1) and (2) can be discretized by a constant step size $\mathbf{\Delta}$ through the zero-order hold rule:
@@ III-F Interpretation of $\mathbf{\Delta}$
[…]
@@ III-H Continuous-Time Prediction
where $v_{next\_time}$ is the interaction item at time next_timenext\_time, $\mathbf{o}\in\mathbb{R}^{D}$ denotes the last item embedding in the output embedding of the last SS4Rec block, and $\mathbf{x_{i}}\in\mathbb{R}^{D}$ denotes the candidate item embedding. In our model, the learned representation contains the information of the next interaction time, i.e., the time when an item needs to be recommended for the user. Thereby, SS4Rec is capable of selecting the most appropriate items at the optimal time for personalized recommendations.
@@ IV-A1 Baselines
  • GRU4Rec [hidasi2015session]: it employs Gated Recurrent Unit (GRU) to model user sequences and encode user interests into a final hidden state for session-based recommendation, where each user’s sequence of clicks is treated as a session.
  • NARM [li2017neural]: a GRU-based method adopts an encoder-decoder architecture to capture both user’s sequential actions and primary intent within a session.
@@
  • SASRec [kang2018self]: this is a classical sequential recommendation method that leverages the Transformer’s self-attention mechanism to process the input sequence.
  • BERT4Rec [sun2019bert4rec]: it utilizes bidirectional self-attention to model user behavior sequences, allowing each item to incorporated contextual information from both preceding and subsequent interactions in the user’s history.
  • TiSASRec [li2020time]: this approach introduces discrete time interval embedding in SASRec, enabling it to effectively capture temporal information.
@@
  • Mamba4Rec [liu2024mamba4rec]: the first Mamba-based sequential recommender, which mainly replaces the self-attention mechanism with the Mamba layer, in contrast to SASRec.
  • RecMamba [yang2024uncovering]: this is a simplified model that uses only the mamba layer, which can be referred to as S6-only model in the Ablation Study.
@@ IV-A2 Datasets
  • Amazon [mcauley2015image]: a series of datasets are gathered from the Amazon e-commerce platform, comprising a substantial volume of product reviews extracted from the Amazon website. We extract the categories ”video games” and ”sports” for subsequent experiments.
  • Movielens [harper2015movielens]: a widely used dataset collected from the Movielens website records user ratings of movies seen. The version (MovieLens-1M) used in our experiments contains 1 million user ratings.
  • KuaiRec [gao2022kuairec]: a real-world dataset is collected from the recommendation logs of the video-sharing mobile app Kuaishou. It contains two interaction matrices, a large matrix and a small matrix, where the small matrix, referred to as the first dataset, contains a fully observed user-item interaction matrix. In this paper, KuaiRec(Large) refers to the large interaction matrix, while KuaiRec(Small) denotes the small one.
@@ IV-A3 Implementation Details
To assess the efficacy of our sequential recommendation model, we follow the standard evaluation protocol used in this field. Specifically, we apply the leave-one-out strategy for testing within the recbole1.0 code framework [zhao2021recbole]. Hit Ratio@K(HR@K), Normalized Discounted Cumulative Gain@K(NDCG@K), and Mean Reciprocal Rank@K(MRR@K) are utilized to evaluate the recommendation performance, with $K=10$ for all metrics. Higher values for these metrics signify better ranking accuracy, which indicates a more effective recommendation model. The specific hyperparameter settings are depicted in Table II, and all experiments are conducted on a single A800 80GB GPU.
@@ IV-C Time-Dependent Prediction
[…]